# Benchmark stars for mean stellar density and surface gravity estimates of solar-type stars

### v1

@@ 3.1.4 HD 209458
Our estimate for the mean stellar density of HD 209458 is based on the value $\tilde{\rho}=0.733\pm 0.008$ from Southworth (2010) based on the analysis of several light curves obtained from space-based instrumentation described in Southworth (2008) and computed assuming a circular orbit. Crossfield et al. derived a value of $e\cos(\omega)=0.00004\pm 0.00033$ based on their observation of the eclipse of HD 209458 b observed at 24 $\mu$m with Spitzer. To account for the possibility of a slightly eccentric orbit, radial velocity measurements from the following sources have been analysed together with this constraint on $e\cos(\omega)$: 1 – Naef et al. (2004, ELODIE); 2 – Naef et al. (2004, CORALIE); 3 – Butler et al. (2006, HIRES). The values of $T_{0}$ and $P$ were fixed at the values measured by Zellem et al. (2014). The details of this analysis are otherwise very similar to those described for CoRoT-1 in Section 3.1.1. The results are given in Table 8. The mean stellar density allowed for the possibility of an eccentric orbit is $\rho_{\star}=0.730^{+0.011}_{-0.014}\,\rho_{\odot}$.
@@ 3.1.6 WASP-4
The estimate for the mean stellar density of WASP-4 is based on the value $\tilde{\rho}=1.230\pm 0.022$ from Southworth (2012) which is taken from the analysis of high-quality ground-based light curves analysed by Southworth et al. (2009) assuming a circular orbit. Beerer et al. derived a value of $e\cos(\omega)=0.00030\pm 0.00086$ based on their observation of the eclipse of WASP 4 b observed at 3.6 and 4.5 $\mu$m with Spitzer. Radial velocity measurements from the following sources have been analysed together with this constraint on $e\cos(\omega)$: 1 – Knutson et al. (2014, HIRES); 2 – Triaud et al. (2010, CORALIE); 3 – Triaud et al. (2010, HARPS). The values of $T_{0}$ and $P$ were fixed at the values measured by Ivshina & Winn (2022). The deviation in the time of mid-transit from this linear ephemeris due to the slow drift in orbital measured by Ivshina & Winn ($\dot{P}=-5.8±1.6$ ms yr-1) is less than 1  minute over 10 years and so was ignored for this analysis. The details of this analysis are otherwise very similar to those described for CoRoT-1 in Section 3.1.1. The results are given in Table 10. The correction for the eccentricity of the orbit is $\psi=0.999\pm 0.012$, so the mean stellar density is found to be $\rho_{\star}=1.225\pm 0.029\,\rho_{\odot}$.
@@ 3.1.8 WASP-43
WASP-43 is a K7 V star that has a hot-Jupiter companion with a very short orbital period (P=0.81347 d). Bernabò et al. (2025) have published a detailed study of the tidal decay and apsidal motion in this system based on the full-phase light curve acquired by JWST with the MIRI instrument between 6.5 and 7.0 $\mu$m, two sectors of TESS data, new and archival radial velocity measurements, and published times of mid-transit and mid-eclipse. This analysis reveals a small but significant eccentricity in the orbit of WASP-43 b ($e=0.00188\pm 0.00035$). There are three values of $a/R_{\star}$ quoted in this study, one from the analysis of the JWST lightcurves and one from each of the two sectors of TESS data. All these values are based on an eccentric orbit fit (L. Bernabò, priv. comm.) and agree well with each other. Using the weighted mean value $a/R_{\star}=4.845\pm 0.028$ the mean stellar density of WASP-43 is found to be $\rho_{\star}=2.300\pm 0.040\,\rho_{\odot}$.
@@ 3.2 Eclipsing binaries with directly measured mass ratios
There are currently only two eclipsing binaries (EBs) with flux ratios $\lesssim$ 0.2 % at optical wavelengths that have published mass ratio values based on direct measurements using near-infrared spectroscopy – EBLM J0113+31 (Maxted et al., 2022a) and EBLM J0608--59 (Maxted et al., 2024). The primary stars of these (EBs) are both useful benchmark stars for stellar density, and also for stellar mass, radius, luminosity, etc. The apparent magnitudes and flux ratio at several optical and near-infrared wavelengths for these EBs are provided in the Maxted et al. (2022a) and Maxted et al. (2024).
@@ 4.2 Test of method for EBLM systems
EBLM J0113+31 and EBLM J0608--59 are useful test cases for the method to derive the mass and mean stellar density of solar-type stars in EBLM systems that is described in Section 3.3. Despite the extreme flux ratio for these EBLM systems ($\approx 0.1$ per  cent at optical wavelengths), extensive high-quality spectroscopy at near-infrared wavelengths has made it possible to directly measure the mass ratio for these binary systems. This makes it possible to measure the mass, radius, and mean stellar density for both stars in these systems without using stellar models or making assumptions about the properties of the primary star, i.e. without using empirical relations such as Equation (6). These model-independent values of the mass and $\log(\rho/\rho_{\odot})$ are given in Table 1 for EBLM J0113+31 and EBLM J0608--59, and can be compared to the values derived using the method described in Section 3.3.
@@ 4.3 Impact of flux from companion in EBLM systems
Table 2 lists the flux ratio measured from the depth of the secondary eclipse for each of the EBLM systems in the sample. The CHEOPS transmission function is similar to the Gaia G band (Fortier et al., 2024; Gaia Collaboration et al., 2016) so the flux ratio measured from CHEOPS light curves is used directly to estimate the flux ratio in the G band, $\ell_{G}$. To estimate the flux ratio in the V band, $\ell_{\rm V}$, from the flux ratio measured from Kepler light curves, $\ell_{\rm Kp}$, I have used version 2019.3.22 of the colour – effective temperature sequence for dwarf stars compiled from Pecaut & Mamajek (2013) downloaded from Eric Mamajek’s web site. [^8] http://www.pas.rochester.edu/~emamajek/EEM_dwarf_UBVIJHK_colors_Teff.txt Linear interpolation in this sequence is used to estimate (B--V) from Teff for the primary star and from the mass for secondary star. Equations (2a), (6a) and (6b) from Brown et al. (2011) are then used to calculate $({\rm V}-{\rm Kp})\approx 0.346({\rm B}-{\rm V})-0.115$ for each star, ignoring that the small difference between the definitions of the Johnson and Tycho B and V bands. Equation (1) from Stassun et al. (2019) is used to calculate the (G--T) colour from the interpolated values of (BP--RP) and then use the interpolated value of to estimate (G--V) to estimate $\ell_{\rm V}$ from the flux ratio measured from TESS light curves, $\ell_{\rm T}$. The luminosity of the primary star, $L_{1}$, is calculated from $T_{\rm eff}$ and the value of R computed from the values of $M_{\star}$ and $\log(\rho_{\star}/\rho_{\odot})$. To estimate the luminosity ratio, $L_{2}/L_{1}$, I use the mass ratio, $M_{2}/M_{1}$, derived as part of the analysis described in Section 3.3 and the value of $M_{\star}$ given in Table 1 to estimate the mass of the companion star, $M_{2}$. Interpolation in the sequence from Pecaut & Mamajek is then used to find the luminosity of the companion, $L_{2}$, based on its mass, except for EBLM J0113+31 and EBLM J0608--59, where the more accurate values of $L_{1}$ and $L_{2}$ from Maxted et al. (2022a) and Maxted et al. (2024) are used to compute the values of $L_{2}/L_{1}$ given in Table 2.
@@
I used bagemass (Maxted et al., 2015) to estimate $\log(\rho_{\star}/\rho_{\odot})$ for EBLM J0113+31 and EBLM J0608--59 from stellar isochrones computed with the garstec stellar evolution code (Weiss & Schlattl, 2008) based on the observed values of $\log L/L_{\odot}$, Teff and [Fe/H] taken from Maxted et al. (2022a) and Maxted et al. (2024). This was done for two cases for each star, one using the actual value of $\log L/L_{\odot}$ for the primary star and again with the total luminosity of the binary system used instead of $\log L/L_{\odot}$. For EBLM J0113+31, adding the flux of the companion changes the stellar density estimate from $\log(\rho_{\star}/\rho_{\odot})=-0.439$ to $\log(\rho_{\star}/\rho_{\odot})=-0.441$, a difference of $-0.4$ per cent. For EBLM J0608--59, the stellar density estimate changes from $\log(\rho_{\star}/\rho_{\odot})=-0.314$ to $\log(\rho_{\star}/\rho_{\odot})=-0.311$, a difference of $+0.6$ per cent. EBLM J0608--59 has one of the largest values of $L_{2}/L_{1}$ for the selected EBLM systems and the value of $L_{2}/L_{1}$ for EBLM J0113+31 is close to the median value for this sample. The bias in $\log(\rho_{\star}/\rho_{\odot})\approx\pm 0.5$ per cent estimated using bagemass is small compared to the typical error on these values (2.3 per cent).
[…]
@@
EBLM J0500--46 0.16 0.12 0.01 0.19
EBLM J0526--34 0.25 0.21 0.04 0.25
EBLM J0540--17 0.14 0.04 0.07 0.01 0.11
EBLM J0608--59 0.28 0.41 0.08 0.44
EBLM J0627--67 0.25 0.15 0.03 0.37
@@
EBLM J1928--38 0.27 0.13 0.05 0.52
EBLM J1934--42 0.18 0.13 0.21 0.04 0.57
EBLM J2040--41 0.15 0.03 0.09 0.01 0.15
@@ 4.4 Example applications
A comparison between the mass, mean stellar density and surface gravity values from Table 1 to the estimates of these quantities obtained from various catalogues is summarised in Table 3. For each quantity I use the combine function described in the appendix to Maxted et al. (2022b) to compute a weighted mean offset allowing for additional scatter $\sigma_{x}$ in the measurements that is added in quadrature to the quoted standard error. The corrections given by equations (8) and (9) has been applied so that all the comparisons are done against values of $\log(\rho_{\star}/\rho_{\odot})$ and $\log g$ that are consistent with the mass estimates given in each catalogue. The standard errors on $\log(\rho_{\star}/\rho_{\odot})$ and $\log g$ from Table 1 are assumed to be unaffected by this small correction. No value of $d\log(\rho_{\star})/dM_{\star}$ is given for EBLM J0113+31 or EBLM J0608--59 because the mass ratio has been measured directly for these two binary systems, so they are not subject to this small additional source of uncertainty.
Version 8.2 of the TESS input catalogue (TIC v8.2, Stassun et al., 2019) provides estimates for the mass, $\log g$ and $\log(\rho_{\star})$ with standard error estimates for 32 of the stars in Table 1. Stassun et al. state that they can estimate masses for stars “on the main sequence or not too far evolved from it. ” but note that this includes “stars that are subgiants; these should be regarded with caution.”. All the stars in Table 1 are well below the upper limit in radius that Stassun et al. use to separate main-sequence stars and subgiants from red giants. Nevertheless, there is a small ($\approx 10$ per cent) but statistically significant offset in mass compared to the benchmark values in Table 1. It is tempting to blame this on stars that have evolved away from the main sequence as there appears to be a weak trend with surface gravity for stars with $\log g\lesssim 4.0$ visible in Fig. 5. However, there is no such trend in $\log(\rho_{\star})$ apparent in the same figure so there may be a confounding variable at play here that is responsible for these offsets. In any case, the standard error estimates for both mass and $\log(\rho_{\star})$ in TIV v8.2 appear to be a reliable estimate of the precision of these estimates.
@@
Table 3: Summary of comparisons between mass, surface gravity and mean stellar density estimates from various catalogues to the values in Table 1. The number of stars from each catalogue found in Table 1, $n$, is given in the first column. Figures in parentheses give the offset in $(\rho_{\star}/\rho_{\odot})$ as a percentage. The statistical significance of the offset for a two-tailed test is given in the column headed $p$. The quantities used to calculate $\log(\rho_{\star}/\rho_{\odot})$ and $M_{\star}$ are noted in the final column using the column names as specified by the catalogue named in the first column.
@@
TIC v8.2 $\log(\rho_{\star}/\rho_{\odot})$ $-0.040\pm 0.018$ $(-9\%)$ 0.01 $0.004\pm 0.011$ (1%) $\log({\tt rho})$
n=32 $M/M_{\odot}$ $-0.061\pm 0.025$ $(-13\%)$ 0.01 $0.004\pm 0.012$ (1%) mass
asPIC 1.1 $\log(\rho_{\star}/\rho_{\odot})$ $-0.078\pm 0.063$ $(-16\%)$ 0.11 $0.320\pm 0.050$ (109%) $\log({\tt Mass})-3\log({\tt Rad})$
n=29 $M/M_{\odot}$ $-0.000\pm 0.038$ $(-0\%)$ 0.49 $0.167\pm 0.034$ (47%) Mass
StarHorse 2 $\log g$ $+0.002\pm 0.045$ $(+1\%)$ 0.48 $0.250\pm 0.034$ (78%) logg50
n=34 $M/M_{\odot}$ $-0.085\pm 0.032$ $(-18\%)$ 0.01 $0.123\pm 0.041$ (33%) mass50
Gaia DR3 $\log(\rho_{\star}/\rho_{\odot})$ $+0.015\pm 0.032$ $(+4\%)$ 0.31 $0.161\pm 0.025$ (45%) $\log({\tt mass\_flame\_spec})$
$-3\log({\tt radius\_flame\_spec})$
n=28 $M/M_{\odot}$ $-0.036\pm 0.019$ $(-8\%)$ 0.03 $0.080\pm 0.018$ (20%) mass_flame_spec
$\log g$ $-0.178\pm 0.083$ $(-34\%)$ 0.01 $0.328\pm 0.080$ (113%) logg_gspspec_ann

### v2

@@ 3.1.4 HD 209458
[…]
@@ 3.1.6 WASP-4
[…]
@@ 3.1.8 WASP-43
[…]
@@ 3.2 Eclipsing binaries with directly measured mass ratios
There are currently only two eclipsing binaries (EBs) with flux ratios $\lesssim$ 0.2 % at optical wavelengths that have published mass ratio values based on direct measurements using near-infrared spectroscopy – EBLM J0113+31 (Maxted et al., 2022a) and EBLM J0608-59 (Maxted et al., 2024). The primary stars of these (EBs) are both useful benchmark stars for stellar density, and also for stellar mass, radius, luminosity, etc. The apparent magnitudes and flux ratio at several optical and near-infrared wavelengths for these EBs are provided in the Maxted et al. (2022a) and Maxted et al. (2024).
@@ 4.2 Test of method for EBLM systems
EBLM J0113+31 and EBLM J0608-59 are useful test cases for the method to derive the mass and mean stellar density of solar-type stars in EBLM systems that is described in Section 3.3. Despite the extreme flux ratio for these EBLM systems ($\approx 0.1$ per  cent at optical wavelengths), extensive high-quality spectroscopy at near-infrared wavelengths has made it possible to directly measure the mass ratio for these binary systems. This makes it possible to measure the mass, radius, and mean stellar density for both stars in these systems without using stellar models or making assumptions about the properties of the primary star, i.e. without using empirical relations such as Equation (6). These model-independent values of the mass and $\log(\rho/\rho_{\odot})$ are given in Table 1 for EBLM J0113+31 and EBLM J0608-59, and can be compared to the values derived using the method described in Section 3.3.
@@ 4.3 Impact of flux from companion in EBLM systems
Table 2 lists the flux ratio measured from the depth of the secondary eclipse for each of the EBLM systems in the sample. The CHEOPS transmission function is similar to the Gaia G band (Fortier et al., 2024; Gaia Collaboration et al., 2016) so the flux ratio measured from CHEOPS light curves is used directly to estimate the flux ratio in the G band, $\ell_{G}$. To estimate the flux ratio in the V band, $\ell_{\rm V}$, from the flux ratio measured from Kepler light curves, $\ell_{\rm Kp}$, I have used version 2019.3.22 of the colour – effective temperature sequence for dwarf stars compiled from Pecaut & Mamajek (2013) downloaded from Eric Mamajek’s web site. [^8] http://www.pas.rochester.edu/~emamajek/EEM_dwarf_UBVIJHK_colors_Teff.txt Linear interpolation in this sequence is used to estimate (B-V) from Teff for the primary star and from the mass for secondary star. Equations (2a), (6a) and (6b) from Brown et al. (2011) are then used to calculate $({\rm V}-{\rm Kp})\approx 0.346({\rm B}-{\rm V})-0.115$ for each star, ignoring that the small difference between the definitions of the Johnson and Tycho B and V bands. Equation (1) from Stassun et al. (2019) is used to calculate the (G-T) colour from the interpolated values of (BP-RP) and then use the interpolated value of to estimate (G-V) to estimate $\ell_{\rm V}$ from the flux ratio measured from TESS light curves, $\ell_{\rm T}$. The luminosity of the primary star, $L_{1}$, is calculated from $T_{\rm eff}$ and the value of R computed from the values of $M_{\star}$ and $\log(\rho_{\star}/\rho_{\odot})$. To estimate the luminosity ratio, $L_{2}/L_{1}$, I use the mass ratio, $M_{2}/M_{1}$, derived as part of the analysis described in Section 3.3 and the value of $M_{\star}$ given in Table 1 to estimate the mass of the companion star, $M_{2}$. Interpolation in the sequence from Pecaut & Mamajek is then used to find the luminosity of the companion, $L_{2}$, based on its mass, except for EBLM J0113+31 and EBLM J0608-59, where the more accurate values of $L_{1}$ and $L_{2}$ from Maxted et al. (2022a) and Maxted et al. (2024) are used to compute the values of $L_{2}/L_{1}$ given in Table 2.
@@
I used bagemass (Maxted et al., 2015) to estimate $\log(\rho_{\star}/\rho_{\odot})$ for EBLM J0113+31 and EBLM J0608-59 from stellar isochrones computed with the garstec stellar evolution code (Weiss & Schlattl, 2008) based on the observed values of $\log L/L_{\odot}$, Teff and [Fe/H] taken from Maxted et al. (2022a) and Maxted et al. (2024). This was done for two cases for each star, one using the actual value of $\log L/L_{\odot}$ for the primary star and again with the total luminosity of the binary system used instead of $\log L/L_{\odot}$. For EBLM J0113+31, adding the flux of the companion changes the stellar density estimate from $\log(\rho_{\star}/\rho_{\odot})=-0.439$ to $\log(\rho_{\star}/\rho_{\odot})=-0.441$, a difference of 0.4-0.4 per cent. For EBLM J0608-59, the stellar density estimate changes from $\log(\rho_{\star}/\rho_{\odot})=-0.314$ to $\log(\rho_{\star}/\rho_{\odot})=-0.311$, a difference of +0.6+0.6 per cent. EBLM J0608-59 has one of the largest values of $L_{2}/L_{1}$ for the selected EBLM systems and the value of $L_{2}/L_{1}$ for EBLM J0113+31 is close to the median value for this sample. The bias in $\log(\rho_{\star}/\rho_{\odot})\approx\pm 0.5$ per cent estimated using bagemass is small compared to the typical error on these values (2.3 per cent).
Table 2: Flux and luminosity ratios for EBLM systems. The flux ratios measrured in the CHEOPS, Kepler and TESS bands are denoted $\ell_{\rm G}$, $\ell_{\it Kepler}$ and $\ell_{\rm T}$, respectively. The flux ratio in the V band $\ell_{\rm V}$ and the luminisity ratio $\frac{L_{2}}{L_{1}}$ have been estimated using the method described in Section 4.3.
@@
EBLM J0500-46 0.16 0.12 0.01 0.19
EBLM J0526-34 0.25 0.21 0.04 0.25
EBLM J0540-17 0.14 0.04 0.07 0.01 0.11
EBLM J0608-59 0.28 0.41 0.08 0.44
EBLM J0627-67 0.25 0.15 0.03 0.37
@@
EBLM J1928-38 0.27 0.13 0.05 0.52
EBLM J1934-42 0.18 0.13 0.21 0.04 0.57
EBLM J2040-41 0.15 0.03 0.09 0.01 0.15
@@ 4.4 Example applications
A comparison between the mass, mean stellar density and surface gravity values from Table 1 to the estimates of these quantities obtained from various catalogues is summarised in Table 3. For each quantity I use the combine function described in the appendix to Maxted et al. (2022b) to compute a weighted mean offset allowing for additional scatter $\sigma_{x}$ in the measurements that is added in quadrature to the quoted standard error. The corrections given by equations (8) and (9) has been applied so that all the comparisons are done against values of $\log(\rho_{\star}/\rho_{\odot})$ and $\log g$ that are consistent with the mass estimates given in each catalogue. The standard errors on $\log(\rho_{\star}/\rho_{\odot})$ and $\log g$ from Table 1 are assumed to be unaffected by this small correction. No value of $d\log(\rho_{\star})/dM_{\star}$ is given for EBLM J0113+31 or EBLM J0608-59 because the mass ratio has been measured directly for these two binary systems, so they are not subject to this small additional source of uncertainty.
Version 8.2 of the TESS input catalogue (TIC v8.2, Stassun et al., 2019) provides estimates for the mass, $\log g$ and $\log(\rho_{\star})$ with standard error estimates for 32 of the stars in Table 1. Stassun et al. state that they can estimate masses for stars “on the main sequence or not too far evolved from it. ” but note that this includes “stars that are subgiants; these should be regarded with caution.”. All the stars in Table 1 are well below the upper limit in radius that Stassun et al. use to separate main-sequence stars and subgiants from red giants. Nevertheless, there is a small ($\approx 10$ per cent) but statistically significant offset in mass compared to the benchmark values in Table 1. It is tempting to blame this on stars that have evolved away from the main sequence as there appears to be a weak trend with surface gravity for stars with $\log g\lesssim 4.0$ visible in Fig. 5. However, there is no such trend in $\log(\rho_{\star})$ apparent in the same figure so there may be a confounding variable at play here that is responsible for these offsets. In any case, the standard error estimates for both mass and $\log(\rho_{\star})$ in TIC v8.2 appear to be a reliable estimate of the precision of these estimates.
@@
Table 3: Summary of comparisons between mass, surface gravity and mean stellar density estimates from various catalogues to the values in Table 1. The number of stars from each catalogue found in Table 1, $n$, is given in the first column. Figures in parentheses give the offset in $(\rho_{\star}/\rho_{\odot})$ as a percentage. The statistical significance of the offset for a two-tailed test is given in the column headed pp. The quantities used to calculate $\log(\rho_{\star}/\rho_{\odot})$ and $M_{\star}$ are noted in the final column using the column names as specified by the catalogue named in the first column.
@@
TIC v8.2 $\log(\rho_{\star}/\rho_{\odot})$ $0.040\pm 0.018$ $(-9\%)$ 0.01 $0.004\pm 0.011$ (1%) $\log({\tt rho})$
n=32 $M/M_{\odot}$ $0.061\pm 0.025$ $(-13\%)$ 0.01 $0.004\pm 0.012$ (1%) mass
asPIC 1.1 $\log(\rho_{\star}/\rho_{\odot})$ $0.078\pm 0.063$ $(-16\%)$ 0.11 $0.320\pm 0.050$ (109%) $\log({\tt Mass})-3\log({\tt Rad})$
n=29 $M/M_{\odot}$ $0.000\pm 0.038$ $(-0\%)$ 0.49 $0.167\pm 0.034$ (47%) Mass
StarHorse 2 $\log g$ $0.031\pm 0.041$ $(-7\%)$ 0.21 $0.226\pm 0.031$ (68%) logg50
n=34 $M/M_{\odot}$ $0.085\pm 0.032$ $(-18\%)$ 0.01 $0.124\pm 0.041$ (33%) mass50
Gaia DR3 $\log(\rho_{\star}/\rho_{\odot})$ $0.015\pm 0.032$ $(+4\%)$ 0.31 $0.161\pm 0.025$ (45%) $\log({\tt mass\_flame\_spec})$
$3\log({\tt radius\_flame\_spec})$
n=28 $M/M_{\odot}$ $0.036\pm 0.019$ $(-8\%)$ 0.03 $0.080\pm 0.018$ (20%) mass_flame_spec
$\log g$ $0.213\pm 0.078$ $(-39\%)$ 0.00 $0.324\pm 0.075$ (111%) logg_gspspec_ann